OpenStudy (anonymous):

Simplify. Please show all of your work -3x^2+4y^2-(5x^2-7y^2-1)

jhonyy9 (jhonyy9):

-3x*2 +4y*2 -(5x*2 -7y*2 -1) = -3x*2 +4y*2 -5x*2 -(-7y*2) -(-)1 so because -(-)1=+1 your equation will be = -3x*2 +4y*2 -5x*2 +7y*2 +1 = -8x*2 +11y*2 +1